# Scheduling problem of aircrafts on a same runway and dual runways

### v3

@@ VII Algorithms
VII-A Algorithm 1
Step 2. Search for the optimal sequence of aircraft $Tcf_{11}$ and $Tcf_{12}$.
Step $i$, $i=4,5,\cdots,n$. Suppose that $\phi_{i-1}=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{i-1}\rangle$ is an optimal sequence for aircraft $Tcf_{11},Tcf_{12},\cdots,Tcf_{1(i-1)}$. Search for the optimal sequences of aircraft $Tcf_{11},Tcf_{12},\cdots,Tcf_{1i}$ based on the obtained theoretical results and the main idea is as follows.
$(i-1)$. Based on the obtained optimal sequence $\phi_{i-1}$, by a series of equivalent transformations, record all possible optimal sequences based on the bases of all the breakpoint-drift equivalent transformation sets and the takeoff-landing/landing-takeoff transitions, and represent their formed set as $\Lambda_{i-1}$ for aircraft $Tcf_{11},Tcf_{12},\cdots,Tcf_{1(i-1)}$. To find the optimal sequences of aircraft $Tcf_{11},Tcf_{12},\cdots,Tcf_{1i}$, we can first insert aircraft $Tcf_{1i}$ between any two adjacent aircraft for each sequence in the set $\Lambda_{i-1}$, denoted by $\phi_{(i-1)k}$, under its time window constraints without changing the orders of the aircraft in $\phi_{(i-1)k}$. Let $f^{i}_{inc}$ be the smallest value of the objective functions $F(\cdot)$ among all the generated new sequences.
For the landing/takeoff scheduling problem, from Theorems 2, 5 and 9, the value of the objective function $F(\cdot)$ for the landing/takeoff scheduling problem is heavily related to the separation times between breakpoint aircraft and their trailing aircraft and the number of breakpoint aircraft. To obtain or record all possible optimal sequences based on the bases of all the breakpoint-drift equivalent transformation sets, emphasis should be laid on the breakpoint aircraft. For the scheduling problem of landing and takeoff aircrafts on a same runway, the number of breakpoint aircrafts is usually small but the number of the takeoff-landing/landing-takeoff transitions might be large, and emphasis should be laid on the takeoff-landing/landing-takeoff transitions.
Therefore, the goal of this step is actually to find all possible combinations of three consecutive aircrafts of different classes in all sequences of the set $\Lambda_{i-1}$ to obtain the smallest value of the objective functions $F(\cdot)$, where each combination corresponds to an time increment when the aircraft $Tcf_{1i}$ is inserted. Starting from this goal, we can reduce some computation amount by giving up some unnecessary operations. For example, since we are concerned about only the time increment when the aircraft $Tcf_{1i}$ is inserted into $\phi_{(i-1)k}$, we need not to repeatedly consider the insertion of aircraft $Tcf_{1i}$ into the same class combination of aircraft and the same separation times, and only focus on the different class combination of aircraft or different separation times, which might significantly reduce the computation amount.
Time increment would be used frequently in our algorithm. To illustrate its role specifically, we give an example. Generate two landing sequences $\phi_{1}=\langle Tcf_{1},Tcf_{4},Tcf_{2},Tcf_{3}\rangle$ and $\phi_{2}=\langle Tcf_{1},Tcf_{2},Tcf_{4},Tcf_{3}\rangle$ by inserting aircraft $Tcf_{4}$ into a sequence $\phi_{0}=\langle Tcf_{1},Tcf_{2},Tcf_{3}\rangle$. The time increment of aircraft $Tcf_{1}$ and $Tcf_{2}$ in $\phi_{1}$ with respect to $\phi_{0}$ is $\Delta_{1}=S_{14}(\phi_{1})+S_{42}(\phi_{1})-S_{12}(\phi_{0})$ and the time increment of aircraft $Tcf_{1}$ and $Tcf_{2}$ in $\phi_{1}$ with respect to $\phi_{0}$ is $\Delta_{2}=S_{24}(\phi_{2})+S_{43}(\phi_{2})-S_{23}(\phi_{0})$. If $S_{23}(\phi_{1})=S_{23}(\phi_{0})$ and $S_{12}(\phi_{2})=S_{12}(\phi_{0})$, then $F(\phi_{2})-F(\phi_{1})=\Delta_{2}-\Delta_{1}$.
[…]
$(i-3)$ According to the elements of the set $F_{inc}$, adjust the aircraft orders in the sequence $\phi_{(i-1)k}$ to generate a new sequence $\bar{\phi}_{(i-1)k}$ which contains a subsequence $\langle Tcf_{p_{1}},Tcf_{k_{1}},Tcf_{1i},Tcf_{j_{1}}\rangle$ such that $\langle p_{1},{k_{1}},{j_{1}}\rangle\in F_{inc}$. Calculate $F(\bar{\phi}_{ik})$ and compare the values of all possible $F(\bar{\phi}_{ik})$ and $f^{i}_{inc}$ so as to find the optimal sequence of aircraft $Tcf_{11},Tcf_{12},\cdots,Tcf_{1i}$.
To generate the subsequence $\langle Tcf_{p_{1}}Tcf_{k_{1}},Tcf_{j_{1}}\rangle$, we can first move the aircraft $\langle Tcf_{p_{1}}Tcf_{k_{1}},Tcf_{j_{1}}\rangle$ to proper place by equivalent transformations including breakpoint-drift equivalent transformations and then use the extraction operation and insertion operation to generate the subsequence $\langle Tcf_{p_{1}}Tcf_{k_{1}},Tcf_{j_{1}}\rangle$ with a focus on the breakpoint aircraft and the landing-takeoff/takeoff-landing transitions on the basis of the optimal sequence.
Since $Y_{p_{1}k_{1}}+Y_{k_{1}(1i)}+Y_{(1i)j_{1}}-Y_{p_{1}j_{1}}<f^{i}_{inc}-F(\phi_{i-1})$, there is no utilizable subsequence $\langle Tcf_{p_{1}},Tcf_{k_{1}},Tcf_{j_{1}}\rangle$ in each $\phi_{(i-1)k}$ for $Tcf_{1i}$ to be inserted between aircraft $Tcf_{k_{1}}$ and $Tcf_{j_{1}}$. So, the generation of a utilizable subsequence $\langle Tcf_{p_{1}}Tcf_{k_{1}},Tcf_{j_{1}}\rangle$ would increase the value of the objective function $F(\cdot)$ for aircraft $Tcf_{11},Tcf_{12},\cdots,Tcf_{1(i-1)}$. Moreover, when $f^{i}_{inc}-F(\phi_{i-1})>T_{0}$, for the landing/takeoff scheduling problem, the optimal insertion of $Tcf_{1(i-1)}$ might usually create a new breakpoint or change an existing breakpoint with some necessary aircraft order adjustments, whereas, for the landing and takeoff scheduling problem on a same runway, the optimal insertion of $Tcf_{1(i-1)}$ might usually create a new breakpoint or change an existing breakpoint with some necessary aircraft order adjustments since the number of breakpoint aircraft is not large. These observations can also be used to reduce the computation amount of Algorithm 1.
Remark 35
For the landing/takeoff scheduling problem, except for the scenarios where aircraft $Tcf_{1i}$ is inserted between a breakpoint aircraft and its trailing aircraft, if each insertion of aircraft $Tcf_{1i}$ generates a new breakpoint aircraft in step $(i-1)$, then aircraft $Tcf_{1i}$ or its leading aircraft in the optimal sequences for aircraft $Tcf_{11},Tcf_{12},\cdots,Tcf_{1i}$ might be a breakpoint aircraft.
Remark 36
Since different classes might correspond to different time increments, we can search for the desired aircraft of proper classes in $F_{inc}$ in ascending order of time increments to minimize the value of the objective function $F(\cdot)$. Moreover, note that the separation times between aircraft of different classes are time-invariant and the number of the classes in $\mathcal{I}$ is not very large. The computation cost of time increments is also not very large and can even be calculated out offline.
Remark 37
It should be noted that in Algorithm 1, when there is no resident-point aircraft, the optimal sequence $\phi_{i}$ satisfies that $F(\phi_{i-1},Sr(\phi_{i-1}))\leq F(\phi_{i},Sr(\phi_{i}))\leq F(\phi_{i-1},Sr(\phi_{i-1}))+Y_{(i-1)i}$ for the landing/takeoff scheduling problem and the landing and takeoff scheduling problem on a same runway, where $T_{0}\leq Y_{(i-1)i}\leq 3T_{0}$. For example, consider a sequence $\phi_{0}=\langle Tcf_{1},Tcf_{2}\rangle$, where $Tcf_{1},Tcf_{2}$ are both landing aircraft and $Y_{12}>T_{D}+D_{T}$. Generate a new sequence $\phi_{1}$ by inserting a takeoff aircraft $Tcf_{3}$ to be between $Tcf_{1}$ and $Tcf_{2}$, where $S_{13}(\phi_{1})=T_{D}$ and $S_{12}(\phi_{1})=S_{12}(\phi_{0})$. It is clear that $F(\phi_{1})-F(\phi_{0})=0$.
Remark 38
Remark 39
VII-B Algorithm 2
Algorithm 2. Suppose that there are totally $n+m$ aircraft composed of $n$ landing aircraft and $m$ takeoff aircraft, denoted by $Tcf_{1},Tcf_{2},\cdots,Tcf_{n+m}$, and there is at least a feasible sequence for them to land/takeoff without conflicts. Let $Tcf^{0}_{1},Tcf^{0}_{2},\cdots,Tcf^{0}_{n}$ denote all the landing aircraft, and $Tcf^{1}_{1},Tcf^{1}_{2},\cdots,Tcf^{1}_{m}$ denote all the takeoff aircraft. The main steps are as follows.
When our algorithm is applied, in each landing/takeoff sequence of each block subsequence, we only consider the case of one breakpoint aircraft and the case of two or more breakpoint aircraft in a block subsequence can be dealt with by combing the operations of extraction and insertion to decrease the value of the objective functions based on the analysis of the case of one breakpoint aircraft for block subsequences. Or, the case of two or more breakpoint aircraft in a block subsequence can be addressed offline as the case of one breakpoint aircraft.
As discussed in Sec. VI, the number of all possible block subsequences of different class combinations is not large according to the standards at the minimum separation time standards at Heathrow Airport and for the RECAT-EU system, when resident-point aircraft are not taken into account.
(2.2) Focus on combinations of different block subsequences, and study the landing/takeoff time increments of any two given consecutive block subsequences, where the last two aircraft of the first block subsequence is the first two aircraft of the second block subsequence. According to the landing/takeoff time increments of two consecutive block subsequences, classify all possible two consecutive subsequences into several sets.
In our simulations for the case of $F(\Phi_{a})\geq F(\Phi_{b})$, the combination of a D-block subsequence and a T-block subsequence is commonly used, which has the form of $\langle DB,TB\rangle$ where $DB$ denotes a D-block subsequence and $TB$ denotes a T-block subsequence, the first two aircraft of $DB$ have the same landing/takeoff time, and the last two aircraft of $TB$ have the same landing/takeoff time. The landing time increment of T-block subsequence in such a combination is a key quantity that needs to be studied for the optimal sequence.
Under Assumption 17, the sequence $\Phi_{c}$ can be divided into a series of block subsequences. Since $F(\Phi^{c}_{LA},Sr_{LA}(\Phi_{c}))=F(\Phi_{a})$, we can use a dynamic programming approach to find the optimal sequence $\Phi_{c}$ based on the sets defined in Step 2. It should be noted that at each step of the dynamic programming approach, we need consider all the possible block subsequences that can decrease the objective function $F(\cdot)$, and if the generation of one block subsequence might disrupt the existing block subsequences, we can first generate the block subsequence and then use it as a basis to generate other block subsequences in an optimal way.
Here, each sequence of $\Lambda_{a}$ is required to be considered. Actually, we usually need to consider only the impact of the local differences of the sequence from other sequences, which might not need a lot of the computation amount.
Step 4. (4.1) Generate a new sequence $\Phi_{c4}$ with $F(\Phi_{c4})\in[F(\Phi_{a}),F(\Phi_{c3})]$ by increasing the separation times between adjacent aircraft in each landing sequence in $\Lambda_{a}$ to form D-block subsequences. Specifically, according to the time increments of block subsequences discussed in Step 2, based on each landing sequence in $\Lambda_{a}$, we first select proper aircraft from $\Phi^{sub}_{c}$ to generate D-block subsequences especially with some or all of the aircraft $\langle Tcf_{n+m-n_{z}+1},Tcf_{n+m-n_{z}+2},\cdots,Tcf_{n+m}\rangle$, denoted by $B_{1},B_{2},\cdots,B_{s}$ for some integer $s>0$. On the basis of these block subsequences, generate a new sequence $\Phi_{c4}$ and minimize $F(\Phi_{c4})$ for all the aircraft.
When increasing the separation times between adjacent aircrafts to form block subsequences, some necessary aircraft order adjustments might be included. For example, suppose that $\phi=\langle Tcf_{1},Tcf_{2},Tcf_{3},Tcf_{4},Tcf_{5}\rangle$ is a landing sequence. When generating block subsequences by increasing the separation times between aircraft $Tcf_{2}$ and $Tcf_{3}$, if the objective function value of the landing sequence can be decreased by exchanging the orders of $Tcf_{3}$ and $Tcf_{5}$, we need to consider the block subsequences based on the sequence $\phi=\langle Tcf_{1},Tcf_{2},Tcf_{5},Tcf_{4},Tcf_{3}\rangle$.
Another approach is to generate a new sequence directly according to block subsequences and the combinations of block subsequences discussed in Step 2 but might have more computation amount.
If $F(\Phi_{c4})\leq F(\Phi_{c3})$, then the global optimal value of the objective function $F(\cdot)$ lies in $[F(\Phi_{a}),F(\Phi_{c4})]$. If $F(\Phi_{c4})>F(\Phi_{c3})$, then the global optimal value of the objective function $F(\cdot)$ still lies in $[F(\Phi_{a}),F(\Phi_{c3})]$.
If the sum of the landing time increments in block subsequences $B_{1},B_{2},\cdots,B_{s}$ is larger than $F(\Phi_{c3})-F(\Phi_{a})$, then $F(\Phi_{c4})>F(\Phi_{c3})$ since the landing sequence $\Phi^{c4}_{LA}$ of $\Phi_{c4}$ might not belong to the set of all possible optimal landing sequences $\Lambda_{a}$. Furthermore, if $F(\Phi^{c4}_{LA},Sr_{LA}(\Phi_{c4}))>F(\Phi_{c3})$, it follows that $F(\Phi_{c4})>F(\Phi_{c3})$.
It should be noted that some or all of the aircraft $Tcf_{n+m-n_{z}+1},Tcf_{n+m-n_{z}+2},\cdots,Tcf_{n+m}$ can be selected to substitute the aircraft of $\Phi^{sub}_{c}$ without changing the intervals between aircraft in $\Phi^{sub}_{c}$.
Remark 40
The main idea of Algorithm 2 is to decompose the aircraft sequence into several block subsequences and fully explore combinations of the block subsequences along an optimal landing/takeoff subsequence to consider the optimization problem (6). Note that the number of all possible block subsequences of different class combinations is not large according to the standards at the minimum separation time standards at Heathrow Airport and for the RECAT-EU system, when resident-point aircraft are not taken into account. The use of block subsequences might significantly reduce the computation amount of the algorithm.
Remark 41
The complexities of Algorithms 1 and 2 are heavily related to the aircraft number, the aircraft classes, and the constraints of the aircraft time windows. But the algorithms essentially exhibit the features of polynomial algorithms and can be applied in actual systems in real time.
VII-C Some necessary results for Algorithms 1 and 2
In Algorithms 1 and 2, the following definitions, lemmas and theorems might be used.
@@ VII-C Some necessary results for Algorithms 1 and 2
From this example, it can also be seen that the bases of all the breakpoint-drift equivalent transformation sets can be interconverted by focusing on the orders of the breakpoint aircrafts and their trailing aircrafts.
@@
Consider a landing/takeoff sequence $\phi=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\rangle$. Suppose that $[f^{0},f^{1}]\subset\bigcap_{i=1}^{n}[f_{i}^{\min},f_{i}^{\max}]$. In the time interval $[f^{0},f^{1}]$, there exists at least 00 to $2$ breakpoint aircraft.

### v5

@@ VII Algorithms
VII-A Some necessary results
In this subsection, we present some necessary definitions, lemmas and theorems for the algorithms.
@@ VII-A Some necessary results
From this example, it can also be seen that the bases of all the breakpoint-drift equivalent transformation sets can be interconverted by focusing on the orders of the breakpoint aircraft and their trailing aircraft.
@@
Proof: This lemma is obvious and its proof is omitted.
@@
Proof: The statement (1) holds naturally from the optimality of the sequence $\Phi_{a}$. Suppose that $j-i>1$, $k-j>1$, and $h-k>1$. From Theorems 1 and 4, the movement of $Tcf_{i}$ from $\Phi_{a}$ to $\Phi_{b}$ and the movement of $Tcf_{j}$ from $\Phi_{b}$ to $\Phi_{c}$ have no effects with each other. Generate a new sequence $\Phi_{d}$ by moving an aircraft $Tcf_{j}$ to be between $Tcf_{h}$ and $Tcf_{h+1}$ in $\Phi_{a}$ under Assumptions 1-10. It follows that $F(\Phi_{c})-F(\Phi_{a})=F(\Phi_{b})-F(\Phi_{a})+F(\Phi_{d})-F(\Phi_{a})$. If $F(\Phi_{b})>F(\Phi_{a})$ and $F(\Phi_{c})=F(\Phi_{a})$, it follows that $F(\Phi_{d})<F(\Phi_{a})$, which is a contradiction. Therefore, $F(\Phi_{b})=F(\Phi_{a})$.
@@
Proof: Note that $f_{1}^{\min}\leq f_{2}^{\min}\leq\cdots\leq f_{i}^{\min}$ and there is at least a sequence for all aircraft to land/take off without conflicts. Without loss of generality, denote $\phi=\langle\phi_{0},\phi_{1}\rangle$, where $\phi_{1}=\langle Tcf_{k_{0}},Tcf_{k_{1}},\cdots,Tcf_{k_{h}}\rangle$ with $k_{0}=i$ for some integer $h>0$. If aircraft $Tcf_{k_{j}}$ is an resident-point aircraft for some integer $0\leq j\leq h$, thus $S_{k_{j-1}k_{j}}(\phi)-Y_{k_{j-1}k_{j}}>0$. Let the landing/takeoff times of the aircraft $Tcf_{k_{j}},Tcf_{k_{j+1}},\cdots,Tcf_{k_{h}}$ be advanced by $S_{k_{j-1}k_{j}}(\phi)-Y_{k_{j-1}k_{j}}>0$. Let $\overline{Sr}(\phi)$ denote the new landing/takeoff times. It is clear that $F(\phi,Sr(\phi))>F(\phi,\overline{Sr}(\phi))$, implying that $(\phi,Sr(\phi))$ is not the optimal solution of $Tcf_{1},Tcf_{2},\cdots,Tcf_{i}$, which is a contradiction.
@@
Consider a landing/takeoff sequence $\phi=\langle Tcf_{1},Tcf_{2},\cdots,Tcf_{n}\rangle$. Suppose that $[f^{0},f^{1}]\subset\bigcap_{i=1}^{n}[f_{i}^{\min},f_{i}^{\max}]$. In the time interval $[f^{0},f^{1}]$, the minimum number of breakpoint aircraft lie in $[0,2]$.
Proof: Since $[f^{0},f^{1}]\subset\bigcap_{i=1}^{n}[f_{i}^{\min},f_{i}^{\max}]$, the orders of aircraft in the time interval $[f^{0},f^{1}]$ can be adjusted arbitrarily. Thus, there might be 00 breakpoint aircraft. Note that the aircraft in the time interval $[t_{0},f^{0})$ and the aircraft in the time interval $[f^{1},+\infty)$ might form class-monotonically-decreasing sequences that finally generate $1$ or $2$ breakpoint aircraft in the time interval $[f^{0},f^{1}]$. Summarizing all the analysis, in the time interval $[f^{0},f^{1}]$, the minimum number of breakpoint aircraft lie in $[0,2]$.